OpenStudy (anonymous):

Two similar cylinders have heights 12 inches and 16 inches. What is the ratio of their surface areas? 1 : 4 9 : 16 3 : 4 16 : 27

OpenStudy (anonymous):

surface area of a cylinder is \[2 \pi r h\] r=radius of cylinder h=height of cylinder

OpenStudy (anonymous):

how do i do that without knowing the radius?

OpenStudy (anonymous):

u have 2 Two similar cylinders that means \[r_{1}=r_{2}\]

OpenStudy (anonymous):

so i just make up a radius?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so i got 226:301

OpenStudy (anonymous):

i used 3 as the radius

OpenStudy (anonymous):

thats right because 226:301=3:4 but u can do it in a better way (surface of cylinder 1)/(surface of cylinder 2)=(2 pi r h1)/(2 pi r h2)=12/16=3/4 since 2 pi r=2 pi r

OpenStudy (anonymous):

yeah i like that way...haha...but thanks :)

OpenStudy (anonymous):

your welcome